# Double Robust high dimensional alpha test for linear factor pricing model

### v1

@@ A.1 Proof of Theorem 1
Similar to the proof of Lemma 2 in Liu et al. (2024), we obtain the following lemma.
LEMMA 3
(Gaussian approximation)Assume Condition (C1)-(C4) with $a_{0}(N)\asymp N^{1-\delta}$ for some positive constant $\delta\leqslant 1/2$ hold. If $\log N=o\left(T^{1/5}\right)$ and $\log T=o\left(N^{1/3\wedge\delta}\right)$, then
$\rho_{T}\left(\mathcal{A}^{\mathrm{re}}\right)=\sup_{A\in\mathcal{A}^{\mathrm{re}}}\left|\mathbb{P}\left\{T^{1/2}\hat{\mathbf{D}}^{-1/2}(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta})\in A\right\}-\mathbb{P}(\boldsymbol{G}\in A)\right|\rightarrow 0,$
as $T\rightarrow\infty$, where $\boldsymbol{G}\sim N\left(0,\zeta_{1}^{-2}\boldsymbol{\Sigma}_{u}\right)$ with $\boldsymbol{\Sigma}_{u}=\mathbb{E}\left({\bf U}_{1}{\bf U}_{1}^{\top}\right)$.
Let $\boldsymbol{Z}=(Z_{1},\dots,Z_{N})\sim N\left(0,E(r_{t}^{2})^{-1}\zeta_{1}^{-2}{\bf R}\right)$. By Condition (C3) and Theorem 2 in Feng et al. (2022a), we have $\mathbb{E}Z_{t}^{2}=\zeta_{1}^{-2}E\left(r_{t}^{2}\right)^{-1}\leq\bar{B}$ and $\mathbb{E}\left[\max_{1\leq j\leq N}Z_{j}\right]\asymp(\sqrt{\log N+\log\log N})$. Let
$\Delta_{0}=\max_{1\leq j,k\leq N}\left|E\left(r_{t}^{2}\right)^{-1}\left(\mathbb{E}{\bf U}_{1}{\bf U}_{1}^{\top}\right)_{j,k}-{\bf R}_{j,k}\right|,$
@@
$\Delta_{0}=\max_{1\leq j,k\leq N}\left|E\left(r_{t}^{2}\right)^{-1}\left(\mathbb{E}{\bf U}_{1}{\bf U}_{1}^{\top}\right)_{j,k}-\mathbf{R}_{j,k}\right|=O\left(N^{-\delta/2}\right).$
@@
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\boldsymbol{Z}\|_{\infty}\leqslant t\right)-\mathbb{P}\left(\|\boldsymbol{G}\|_{\infty}\leqslant t\right)\right|\leqslant C^{\prime}T^{-1/3}(1\vee\log(NT))^{2/3}\rightarrow 0$
@@
$\mathbb{P}\left(E(r_{t}^{2})\zeta_{1}^{2}\max_{1\leq i\leq N}Z_{i}^{2}-2\log N+\log\log N\leq x\right)\rightarrow G(x)=\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-x/2}\right\},$
a cdf of the Gumbel distribution, as $N\rightarrow\infty$. Thus, according to Theorem 1 in Liu et al. (2024).
$\left|\mathbb{P}\left(T||\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}||^{2}_{\infty}\zeta-2\log N+\log\log N\leq x\right)-G(x)\right|$
$\leq$ $\left|\mathbb{P}\left(\zeta T||\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}||^{2}_{\infty}-2\log N+\log\log N\leq x\right)-\mathbb{P}\left(\zeta^{2}\max_{1\leq i\leq N}Z_{i}^{2}-2\log N+\log\log N\leq x\right)\right|$
$+\left|\mathbb{P}\left(\zeta^{2}\max_{1\leq i\leq N}Z_{i}^{2}-2\log N+\log\log N\leq x\right)-G(x)\right|+o(1)\rightarrow 0,$
@@
Due to (1) and $\|\hat{\boldsymbol{\theta}}\|=O_{p}(\zeta_{1}^{-1}T^{-1/2})$, we have $r_{t}^{-2}\left\|\hat{\mathbf{D}}^{-1/2}(\hat{\boldsymbol{\theta}}+\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t})\right\|^{2}=O_{p}\left(T^{-1}\right)=o_{p}(1)$ and by the Cauchy inequality, the other parts are also $o_{p}(1)$. So,
@@
$\left(T^{-1}\sum_{t=1}^{T}\left\|\mathbf{D}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t}\right\|^{2}\right)\mathop{\rightarrow}\limits^{p}E(r_{t}^{2})\text{ and }\left(T^{-1}\sum_{t=1}^{T}\left\|\mathbf{D}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t}\right\|^{-1}\right)\mathop{\rightarrow}\limits^{p}\zeta_{1}.$
Obviously, $\hbox{var}\left(T^{-1}\zeta_{1}^{-1}\sum_{t=1}^{T}r_{t}^{-1}\right)=O\left(T^{-1}\right)$ and $\hbox{var}\left(T^{-1}\sum_{t=1}^{T}r_{t}^{2}\right)=O\left\{T^{-1}(Er_{t}^{2})^{2}\right\}$. Finally, the proof is completed due to Slutsky’s theorem. $\Box$
@@ A.3 Proof of Theorem 2
Without loss of generality, we assume that $\|\boldsymbol{\alpha}\|_{\infty}=O(\sqrt{\log N/T})$ and ${\bf Z}_{t}=\boldsymbol{\varepsilon}_{\cdot t}+(1-\eta_{t})\boldsymbol{\alpha}-\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}$, where $\eta_{t}=\boldsymbol{1}_{T}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}$ and $\boldsymbol{V}_{t}={\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}$. Define $\eta=E(\eta_{t})$, $\hat{{\bf U}}_{t}=U(\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t})$ and $\hat{r}_{t}=\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t}\|$ The estimator $\tilde{\boldsymbol{\theta}}=\hat{\boldsymbol{\theta}}-(1-\eta)\boldsymbol{\alpha}$ satisfies $\sum_{t=1}^{T}U(\hat{{\bf D}}^{-1/2}({\bf Z}_{t}-(1-\eta)\boldsymbol{\alpha}-{\boldsymbol{\theta}}))=\mathbf{0}$, which is equivalent to
@@
×$\times(1+\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}\|^{2}+\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}\|^{2}-2\hat{r}^{-2}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\tilde{\boldsymbol{\theta}}$
$\quad\quad-2\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}+\hat{r}_{t}^{-2}\|\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}\|^{2}-2\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}$
$\quad\quad-2\hat{r}^{-2}_{t}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}+2\hat{r}_{t}^{-2}\boldsymbol{V}^{\top}_{t}\boldsymbol{\varepsilon}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}+2\hat{r}_{t}^{-2}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha})^{-1/2}=\mathbf{0}$
@@
×$\times(1-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}\|^{2}-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}\|^{2}+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\tilde{\boldsymbol{\theta}}$
$\quad\quad+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}+\frac{1}{2}\hat{r}_{t}^{-2}\|\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}\|^{2}-\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}$
@@
$\quad\quad+\hat{r}^{-2}_{t}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}+\hat{r}_{t}^{-2}\boldsymbol{V}^{\top}_{t}\boldsymbol{\varepsilon}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}+\hat{r}_{t}^{-2}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}+\delta_{3t})=\mathbf{0},$

### v2

@@ A.1 Proof of Theorem 1
Let $L_{N,T}=T^{-1/4}\log^{1/2}(NT)+N^{-(1/6\wedge\delta/2)}\log^{1/2}(NT)+T^{-1/2}(\log N)^{1/2}\log^{1/2}(NT)$. Then for any sequence $\eta_{T}\rightarrow\infty$ and any $t\in\mathbb{R}^{p}$,
$P\left(T^{1/2}\hat{\mathbf{D}}^{-1/2}(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta})\leq t\right)=$ $P\left(T^{-1/2}\zeta_{1}^{-1}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st})\boldsymbol{U}_{t}+C_{T}\leq t\right)$
$\leq$ $P\left(T^{-1/2}\zeta_{1}^{-1}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st})\boldsymbol{U}_{t}\leq t+\eta_{T}L_{N,T}\right)$
$+P\left(\left\|C_{T}\right\|_{\infty}>\eta_{T}L_{N,T}\right).$
When $r_{1},\dots,r_{N},{\boldsymbol{f}}_{1}\dots,{\boldsymbol{f}}_{T}$ are given, according to Lemma A4 in Cheng et al. (2023) and the Gaussian approximation for independent partial sums in Koike (2021), let
$\boldsymbol{G}\sim N\left(0,\zeta_{1}^{-2}T^{-1}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st})^{2}\mathbb{E}\left(\boldsymbol{U}_{1}\boldsymbol{U}_{1}^{\top}\right)\right),$
we can derive that the the conditional probability is
$P_{1}\left(T^{1/2}\zeta_{1}^{-1}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st})\boldsymbol{U}_{t}\leq t+\eta_{T}L_{N,T}\right)$
$\leq$ $P_{1}\left(\boldsymbol{G}\leq t+\eta_{T}L_{N,T}\right)+O\left(\left\{T^{-1}\log^{5}(NT)\right\}^{1/6}\right)$
$\leq$ $P_{1}(\boldsymbol{G}\leq t)+O\left\{\eta_{T}L_{N,T}\log^{1/2}(N)\right\}+O\left(\left\{T^{-1}\log^{5}(NT)\right\}^{1/6}\right),$
where the second inequality holds due to Lemma 12 in Liu et al. (2024). Note that
$T^{-1}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st})^{2}$
$=$ $1-\frac{2}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st}+\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}r_{s}^{-2}r_{t}^{2}V_{st}^{2}$
$+$ $\frac{1}{T}\sum_{t=1}^{T}\underset{s_{1}\neq s_{2}}{\sum^{T}\sum^{T}}r_{s_{1}}^{-1}r_{s_{2}}^{-1}r_{t}^{2}V_{s_{1}t}V_{s_{2}t}\mathop{\rightarrow}\limits^{p}1-2\eta E(r_{t}^{-1})E(r_{t})+\eta E(r_{s}^{-2})E(r_{t}^{2}),$
where $T^{-1}{\bf 1}_{T}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\bf F}^{\top}{\bf 1}_{T}\mathop{\rightarrow}\limits^{p}\eta$. Thus, let $1-2\eta E(r_{t}^{-1})E(r_{t})+\eta E(r_{s}^{-2})E(r_{t}^{2})=\eta_{\omega}$ and $\boldsymbol{G}^{{}^{\prime}}\sim N\left(0,\zeta_{1}^{-2}\eta_{\omega}\mathbb{E}\left(\boldsymbol{U}_{1}\boldsymbol{U}_{1}^{\top}\right)\right),$ we have
$P\left(T^{1/2}\hat{\mathbf{D}}^{-1/2}(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta})\leq t\right)\leq$ $P(\boldsymbol{G}^{{}^{\prime}}\leq t)+O\left\{\eta_{T}L_{N,T}\log^{1/2}(N)\right\}+O\left(\left\{T^{-1}\log^{5}(NT)\right\}^{1/6}\right)$
$+P\left(\left|C_{T}\right|_{\infty}>\eta_{T}L_{N,T}\right)+o(1).$
On the other hand, we have
$P\left(T^{1/2}\hat{\mathbf{D}}^{-1/2}(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta})\leq t\right)\geq$ $P(\boldsymbol{G}^{{}^{\prime}}\leq t)-O\left\{\eta_{T}L_{N,T}\log^{1/2}(N)\right\}-O\left(\left\{T^{-1}\log^{5}(NT)\right\}^{1/6}\right)$
$-P\left(\left\|C_{T}\right\|_{\infty}>\eta_{T}L_{N,T}\right)-o(1),$
where $P\left(\left\|C_{T}\right\|_{\infty}>\eta_{T}L_{N,T}\right)\rightarrow 0$ as $T\rightarrow\infty$. Then we have that, if $\log N=o\left(T^{1/5}\right)$ and $\log T=o\left(N^{1/3\wedge\delta}\right)$,
$\sup_{t\in\mathbb{R}^{p}}\left|P\left(T^{1/2}\hat{\mathbf{D}}^{-1/2}(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta})\leq t\right)-P(\boldsymbol{G}^{{}^{\prime}}\leq t)\right|\rightarrow 0.$
Further,
$\rho_{T}\left(\mathcal{A}^{re}\right)=\sup_{A\in\mathcal{A}^{re}}\left|P\left(T^{1/2}\hat{\mathbf{D}}^{-1/2}(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta})\in A\right)-P(\boldsymbol{G}^{{}^{\prime}}\in A)\right|\rightarrow 0,$
by the Corollary 5.1 in chernozhukov2017central. Let $\boldsymbol{Z}=(Z_{1},\dots,Z_{N})\sim N\left(0,N^{-1}\zeta_{1}^{-2}\eta_{\omega}{\bf R}\right)$. By Condition (C3) and Theorem 2 in Feng et al. (2022a), we have $\mathbb{E}Z_{j}^{2}=\zeta_{1}^{-2}N^{-1}\eta_{\omega}\leq\bar{B}$ and $\mathbb{E}\left[\max_{1\leq j\leq N}Z_{j}\right]\asymp(\sqrt{\log N+\log\log N})$. Let
$\Delta_{0}=\max_{1\leq j,k\leq N}\left|N\left(\mathbb{E}{\bf U}_{1}{\bf U}_{1}^{\top}\right)_{j,k}-{\bf R}_{j,k}\right|,$
@@
$\Delta_{0}=\max_{1\leq j,k\leq N}\left|N\left(\mathbb{E}{\bf U}_{1}{\bf U}_{1}^{\top}\right)_{j,k}-\mathbf{R}_{j,k}\right|=O\left(N^{-\delta/2}\right).$
@@
$\sup_{t\in\mathbb{R}}\left|P\left(\|\boldsymbol{Z}\|_{\infty}\leqslant t\right)-P\left(\|\boldsymbol{G}^{{}^{\prime}}\|_{\infty}\leqslant t\right)\right|\leqslant C^{\prime}T^{-1/3}(1\vee\log(NT))^{2/3}\rightarrow 0.$
@@
$P\left(N\zeta_{1}^{2}\eta_{\omega}^{-1}\max_{1\leq i\leq N}Z_{i}^{2}-2\log N+\log\log N\leq x\right)\rightarrow G(x)=\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-x/2}\right\},$
a cdf of the Gumbel distribution, as $N\rightarrow\infty$. Thus, according to Theorem 2 in Feng et al. (2022a).
$\left|P\left(T||\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}||^{2}_{\infty}\zeta-2\log N+\log\log N\leq x\right)-G(x)\right|$
$\leq$ $\left|P\left(\zeta T||\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}||^{2}_{\infty}-2\log N+\log\log N\leq x\right)-P\left(\zeta^{2}\max_{1\leq i\leq N}Z_{i}^{2}-2\log N+\log\log N\leq x\right)\right|$
$+\left|P\left(\zeta^{2}\max_{1\leq i\leq N}Z_{i}^{2}-2\log N+\log\log N\leq x\right)-G(x)\right|+o(1)\rightarrow 0,$
@@
Due to (A.1) and $\|\hat{\boldsymbol{\theta}}\|=O_{p}(\zeta_{1}^{-1}T^{-1/2})$, we have $r_{t}^{-2}\left\|\hat{\mathbf{D}}^{-1/2}(\hat{\boldsymbol{\theta}}+\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t})\right\|^{2}=O_{p}\left(T^{-1}\right)=o_{p}(1)$ and by the Cauchy inequality, the other parts are also $o_{p}(1)$. So,
@@
$\left(T^{-1}\sum_{t=1}^{T}\left\|\mathbf{D}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t}\right\|^{2}\right)=E(r_{t}^{2})\{1+o_{p}(1)\}\text{ and }\zeta_{1}^{-1}\left(T^{-1}\sum_{t=1}^{T}\left\|\mathbf{D}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t}\right\|^{-1}\right)\mathop{\rightarrow}\limits^{p}1.$
Obviously, $\hbox{var}\left(T^{-1}\zeta_{1}^{-1}\sum_{t=1}^{T}r_{t}^{-1}\right)=O\left(T^{-1}\right)$ and $\hbox{var}\left(T^{-1}\sum_{t=1}^{T}r_{t}^{2}\right)=O\left\{T^{-1}N^{2}\right\}$. Similarly, we can prove that
$\left(T^{-1}\sum_{t=1}^{T}\left\|\mathbf{D}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t}\right\|\right)=E(r_{t})+O_{p}(T^{-1/2}N^{1/2})\text{ and }\zeta_{2}^{-1}\left(T^{-1}\sum_{t=1}^{T}\left\|\mathbf{D}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t}\right\|^{-2}\right)\mathop{\rightarrow}\limits^{p}1.$
Similarly, we can derive that $\hbox{var}\big{(}T^{-1}\sum_{t=1}^{T}\zeta_{2}^{-1}r_{t}^{-2}\big{)}=O\big{(}T^{-1}\big{)}$, $\hbox{var}\big{(}T^{-1}\sum_{t=1}^{T}r_{t}\big{)}=O\big{(}T^{-1}N\big{)}$ and $T^{-1}\omega_{T}\mathop{\rightarrow}\limits^{p}1-\eta.$ Finally, the proof is completed due to Slutsky’s theorem.
$\Box$
@@ A.3 Proof of Theorem 2
Without loss of generality, we assume that $\|\boldsymbol{\alpha}\|_{\infty}=O(\sqrt{\log N/T})$ and ${\bf Z}_{t}=\boldsymbol{\varepsilon}_{\cdot t}+(1-\eta_{t})\boldsymbol{\alpha}-\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}$, where $\eta_{t}=\boldsymbol{1}_{T}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}$ and $\boldsymbol{V}_{t}={\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}$. Recall that $T^{-1}\omega_{T}\mathop{\rightarrow}\limits^{p}1-\eta$, obviously we have $\eta=E(\eta_{t})$. Define $\omega=1-\eta$, $\hat{{\bf U}}_{t}=U(\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t})$ and $\hat{r}_{t}=\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t}\|$. The estimator $\tilde{\boldsymbol{\theta}}=\hat{\boldsymbol{\theta}}-\omega\boldsymbol{\alpha}$ satisfies $\sum_{t=1}^{T}U(\hat{{\bf D}}^{-1/2}({\bf Z}_{t}-\omega\boldsymbol{\alpha}-{\boldsymbol{\theta}}))=\mathbf{0}$, which is equivalent to
@@
×$\times(1+\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}\|^{2}+\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}\|^{2}-2\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}$
$\quad\quad-2\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}+\hat{r}_{t}^{-2}\|\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}\|^{2}-2\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}$
$\quad\quad+2\hat{r}^{-2}_{t}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}+2\hat{r}_{t}^{-2}\boldsymbol{V}^{\top}_{t}\boldsymbol{\varepsilon}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}+2\hat{r}_{t}^{-2}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha})^{-1/2}=\mathbf{0}.$
@@
×$\times(1-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}\|^{2}-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}\|^{2}+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}$
$\quad\quad+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}-\frac{1}{2}\hat{r}_{t}^{-2}\|\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}\|^{2}+\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}$
@@
$\quad\quad-\hat{r}^{-2}_{t}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}-\hat{r}_{t}^{-2}\boldsymbol{V}^{\top}_{t}\boldsymbol{\varepsilon}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}-\hat{r}_{t}^{-2}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}+\delta_{3t})=\mathbf{0},$